Only 27% of U.S adults get enough leisure time to exercise to achieve cardiovascular fitness. Choose 3 adults at random. Find the probability that A: All 3 get enough daily exercise B: At least 1 of the 3 gets enough exercise.


1
Expert's answer
2021-02-26T00:57:53-0500

Let us consider U.S adults get enough time to exercise as a success.

Then probability of success is "\frac{27}{100}" and probability of failure is "\frac{73}{100}".

"\therefore p=\frac{27}{100}" and "q=\frac{73}{100}".

Now we know that for binomial distribution probability of getting "r" success from "n" trial is "={^n}C_r.p^r.q^{(n-r)}"

(A) Here "n=3,r=3"

Therefore probability that all "3" get daily enough exercise is "={^3}C_3.(\frac{27}{100})^3.(\frac{73}{100})^{(3-3)}=(\frac{27}{100})^3" .

(B) First we find out of "3" students no one get enough time to exercise.

Therefore here "n=3,r=0"

So probability that no one get enough exercise is "={^3}C_0.(\frac{27}{100})^0.(\frac{73}{100})^{(3-0)}=(\frac{73}{100})^3" .

Therefore the required probability of getting atleast "1" of the "3" gets enough exercise is "=[1-(\frac{73}{100})^3]" "=0.61"